# Wasserstein–Aitchison GAN for angular measures of multivariate extremes

### v1

@@ Results.
Table 1: Dependence and extremes scores for different dependence $\tau$ and dimensions $d$. For each scenario and each score, the best score (i.e., the lowest) is indicated in bold.
@@
These results show that WA-GAN is competitive with other methods, from the tail dependence perspective as well as from the quality of the generated extremes. The advantage of WA-GAN seems to increase as the dimension $d$ increases. Large values of $d$ is often at the center of interest when generative models for the dependence are used. Note also that it may be surprising at first glance that the scores become better when $d$ increases, but this is probably due to the dependence structure of the data generating process being the same for all pairs of variables, making it easier to learn it when the dimension of the space increases.
@@
Some additional comments can be made for each method. For WA-GAN, the penalization coefficient $\rho$ in (19) is typically selected to be large among the possible values ($\rho=1$ or $\rho=3$) showing that it is a good idea to enforce marginal constraints (3) in the generative model. For HTGAN, the obtained values of the hyperparameters match the discussion on pages 13–15 in the paper [18] proposing this method. In particular, we get big batch sizes, big latent space dimensions, and low values of the learning rate included in the range recommendend by the authors. For GPGAN, the training phase is quicker since, as illustrated on Figure 2, we have a lower training size in this case as this method is trained on points with $|\bm{V}|_{\infty}>n/k$ while WA-GAN is trained on points with $|\bm{V}|_{1}>n/k$, which is less restrictive. This can be an asset if we have enough data, but can harm GPGAN’s performance if $d$ is large and $n$ is not too big.
@@
As an illustration of the performances of the different methods to accurately reproduce the tail dependence of the data, we plot the empirical bivariate and trivariate extremal coefficients ($|J|=2$ and $|J|=3$ in (21)) of the test set against the ones of the training set, and against those of all the generative methods for $d\in\{10,50\}$ and $\tau=0.5$. In this setting, we know that the true values of these coefficients are $\sqrt{|J|}$. This is displayed n Figure 3 for $d\in\{10,50\}$. There is no clear visual difference between WA-GAN and HTGAN, both performing well at representing the coefficients of the test set. GPGAN, however, fails to do so, and even more when the dimension increases. When $d$ gets larger, the empirical extremal coefficients are below their true values, even in the test set. The reason is that the empirical estimator of $\Phi$ in (21) becomes less accurate as $d$ increases [9].
@@
Figure 3: Top: $d=10$; bottom: $d=50$. – Extremal coefficients of order $k=2$ (blue) and $k=3$ (red). Solid black line is the diagonal. Dashed blue and red lines are the true values of the coefficients for $k=2$ and $k=3$ respectively.
@@
Next, to illustrate the accuracy of generated extremes by each method, we plot the projection of each generated sample on the first two margins and compare it to the first two margins of the extremes in the test set. We take again $\tau=0.5$. Results for are displayed on Figure 4. WA-GAN seems to perform well, whatever the dimension. The results of HTGAN are less satisfying as it is clear from the figure that the angular measure associated with this method is discrete (this corresponds to the “directions” that we observe in the generated extremes) and this does not match the true angular measure which is logistic. This is theoretically justified by Proposition 2.10 in [18] introducing this method, and already visible from their Figure 2 on page 14. GPGAN also performs well even though its performance decreases in comparison to WA-GAN as the dimension increases; see Table 1.
@@
Figure 4: Top: $d=10$; bottom: $d=50$. – First two margins of extremes in the test set and the different generative methods.
@@ 4.3 Financial data: daily returns of industry portfolios
[…]
@@
To sample extremes from the joint distribution, we first compute the maximum likelihood estimators of the univariate GPD parameters for each margin. Boxplots illustrating the values of those parameters are reported in Figure 6. The estimated shape values clearly indicates the presence of heavy tails in the marginal distributions. The GPD qq-plots of the marginal fits are to be found on Figures 9, 10 and 11 in Appendix E. They indicate a relatively good fit, even though some extreme quantiles are larger than the ones associated with the GPD model.
Figure 6: Boxplots of estimated marginal GPD parameters
For comparison, we start by computing the test scores for WA-GAN, HTGAN and GPGAN. The training and validation procedures are exactly the same as described in Section 4.2. Here, we train on $n_{\text{train}}=7\,000$, we validate on $n_{\text{val}}=3\,000$ and compute the scores on $n_{\text{test}}=6\,694$ observations. We again consider $k=\sqrt{n_{\text{train}}}$ for WA-GAN. For HTGAN, we do not compute the scores on extremes because the marginal distributions are not know, see Appendix D. Table 2 shows that WA-GAN seems to be the best at capturing the dependence between extremes in the data while GPGAN has a better $2$-Wasserstein distance score.
@@
As already pointed out in the analysis of [25], the extreme losses of those portfolios can be clustered into different kind of industries. For example, the tobacco industry exhibits asymptotic independence to all other categories (e.g., to the textile industries) while the extreme losses of the business and IT-related industries are dependent. We illustrate the fact that WA-GAN also captures those phenomena in Figure 7 by displaying the two-dimensional projections of the estimated angular measure on those margins. We see that the generated “angles” associated with tobacco/textile portfolios are much more concentrated around the axes than for business/IT, showing that WA-GAN captures this complex tail dependence.
Figure 7: Two-dimensional projections of the generated “angles” from $\Phi$ by WA-GAN.
As a further illustration we consider a portfolio obtained by combining the mines industry portfolio with the coal industry portfolio (with equal weight). It is intuitive and has been shown in [10, 25] that the mines and coal industries exhibit asymptotic dependence. If one wants to sample the losses associated to the portfolio we propose, this tail dependence should be taken into account. Figure 8 represents the densities of simulated values of the extreme losses of the combined portfolio (i.e., both marginal losses exceeds a large threshold) based on WA-GAN (in black) and based on independent simulations from the marginal GPDs (in blue). Not taking the tail dependence into account leads to underestimation of the risk associated with this portfolio.
Figure 8: Densities of extreme losses in the combined portfolio of mines and coals industries based on WA-GAN (black) and independently simulated margins (blue).
@@ 5 Conclusions
The method is tried out on simulated extreme value data sets of dimensions $d=10$, $20$ and $50$ with 10 0001000010\,00010 000 observations in the training set and 5 00050005\,0005 000 observations in the validation set. Plots and two performance measures which (i) score the dependence between extremes using extremal coefficients and (ii) score the quality of the simulated values on the real scale using the 2-Wasserstein distance show that WA-GAN performs well. Further, in Table 1 WA-GAN is shown to have better values of these performance scores than the HTGAN and GPGAN methods.
@@
The code is available from the first author and will be made publicly available in a software repository. The data underlying the analysis in Section 4.3 is publicly available from the Kenneth R. French data library https://mba.tuck.dartmouth.edu/pages/faculty/ken.french/data_library.html.
@@ Appendix A Proof of Proposition 1
The exponent measure $\nu$ is connected to the angular measure $\Phi$ introduced in (2). More precisely [7, Chapter 8], for any bounded, measurable function $f:\mathbb{E}\rightarrow\mathbb{R}$ that is zero in a neighborhood of the origin, we have
@@
where $Y$ is unit-Pareto distributed and independent of $\bm{\Theta}\sim\Phi$. Whence, by (26),
@@
From the local uniformity of the convergence in point (ii) of Remark 1, guaranteed by Assumption 3, and the fact that $\mathsf{P}(Y_{j}>0)=1$ for any $j\in\{1,\ldots,d\}$ by the previous computations and Assumption 2, the extended continuous mapping theorem [43, Theorem 1.11.1] applies, yielding
@@
This concludes the proof of (7) and (8). Equation (9) is a direct consequence of (7) by noting that
@@ Appendix B The Aitchison simplex
The open simplex ${\kern 0.0pt\Delta}^{\mathrm{o}}_{d-1}$ is equipped with an inner product space structure with the following operations [1]: for $\bm{v}=(v_{1},\ldots,v_{d}),\bm{w}=(w_{1},\ldots,w_{d})\in{\kern 0.0pt\Delta}^{\mathrm{o}}_{d-1}$ and $\alpha\in\mathbb{R}$, define the following operations:
  • Addition:
    $\bm{v}\oplus\bm{w}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\left(\frac{v_{j}w_{j}}{\sum_{i=1}^{d}v_{i}w_{i}}\right)_{j=1}^{d}$
  • Scalar multiplication:
    $\alpha\odot\bm{v}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\left(\frac{v_{j}^{\alpha}}{\sum_{i=1}^{d}v_{i}^{\alpha}}\right)_{j=1}^{d}$
  • (Aitchison) inner product:
    $\left\langle\bm{v},\bm{w}\right\rangle_{A}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\sum_{i=1}^{d}\log\left(\frac{v_{i}}{g(\bm{v})}\right)\log\left(\frac{w_{i}}{g(\bm{w})}\right),$
    where $g:\bm{u}\in{\kern 0.0pt\Delta}^{\mathrm{o}}_{d-1}\mapsto g(\bm{u})\stackrel{{\scriptstyle\mathrm{def}}}{{=}}(\prod_{i=1}^{d}u_{i})^{1/d}$ is the geometric mean.
It is an easy exercise to verify that $({\kern 0.0pt\Delta}^{\mathrm{o}}_{d-1},\oplus,\odot)$ forms a real vector space with zero element $\bm{0}_{A}=\bm{1}_{d}/d\in{\kern 0.0pt\Delta}^{\mathrm{o}}_{d-1}$, where $\bm{1}_{d}=(1,\ldots,1)\in\mathbb{R}^{d}$, and that $\left\langle\cdot,\cdot\right\rangle_{A}$ defines an inner product on it. The open simplex equipped with this structure is often called the Aitchison simplex.
@@
[…]
@@
$\left\{\bm{x}_{i}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}(0,\ldots,0,1,-1,0,\ldots,0)\in\mathbb{H}:i=1,\ldots,d-1\right\},$

### v2

@@ Results.
Table 1: Logistic dependence structure. Dependence and extremes scores for different dependence $\tau$ and dimensions $d$. For each scenario and each score, the best score (i.e., the lowest) is indicated in bold.
@@
These results show that WA-GAN is competitive with other methods, from the tail dependence perspective as well as from the quality of the generated extremes. The advantage of WA-GAN seems to increase as the dimension $d$ increases. Large values of $d$ is often at the center of interest when generative models for the dependence are used. Note also that it may seem surprising at first that the Wasserstein scores improve as $d$ increases. This can be explained by several factors. One possible explanation in the present setting is that the dependence structure of the data-generating process is identical across all pairs of variables, which makes it easier to learn as the dimensionality of the space increases. Consequently, to ensure that the comparison remains meaningful, the main interest lies in comparing the values for a given dimension.
@@
Some additional comments can be made for each method. For WA-GAN, the penalization coefficient $\rho$ in (19) is typically selected to be large among the possible values ($\rho=1$ or $\rho=3$) showing that it is a good idea to enforce marginal constraints (3) in the generative model. For HTGAN, the obtained values of the hyperparameters match the discussion on pages 13–15 in the paper [Girard-et-al] proposing this method. In particular, we get big batch sizes, big latent space dimensions, and low values of the learning rate included in the range recommendend by the authors. For GPGAN, the training phase is quicker since, as illustrated on Figure 2, we have a lower training size in this case as this method is trained on points with $|\bm{V}|_{\infty}>n/k$ while WA-GAN is trained on points with $|\bm{V}|_{1}>n/k$, which is less restrictive. This can be an asset if we have enough data, but can harm GPGAN’s performance if $d$ is large and $n$ is not too big.
@@
As an illustration of the performances of the different methods to accurately reproduce the tail dependence of the data, we plot the empirical bivariate and trivariate extremal coefficients ($|J|=2$ and $|J|=3$ in (21)) of the test set against the ones of the training set, and against those of all the generative methods for $d\in\{10,50\}$ and $\tau=0.5$. In this setting, we know that the true values of these coefficients are $\sqrt{|J|}$. This is displayed n Figure 3 for $d\in\{10,50\}$. There is no clear visual difference between WA-GAN and HTGAN, both performing well at representing the coefficients of the test set. GPGAN, however, fails to do so, and even more when the dimension increases. When $d$ gets larger, the empirical extremal coefficients are below their true values, even in the test set. The reason is that the empirical estimator of $\Phi$ in (21) becomes less accurate as $d$ increases [clemencon2023concentration].
@@
Figure 3: Logistic dependence structure. Left: $d=10$; right: $d=50$. – Extremal coefficients of order $k=2$ (blue) and $k=3$ (red). Solid black line is the diagonal. Dashed blue and red lines are the true values of the coefficients for $k=2$ and $k=3$ respectively.
@@
Next, to illustrate the accuracy of generated extremes by each method, we plot the projection of each generated sample on the first two margins and compare it to the first two margins of the extremes in the test set. We take again $\tau=0.5$. Results for are displayed on Figure 4. WA-GAN seems to perform well, whatever the dimension. The results of HTGAN are less satisfying as it is clear from the figure that the angular measure associated with this method is discrete (this corresponds to the “directions” that we observe in the generated extremes) and this does not match the true angular measure which is logistic. This is theoretically justified by Proposition 2.10 in [Girard-et-al] introducing this method, and already visible from their Figure 2 on page 14. GPGAN also performs well even though its performance decreases in comparison to WA-GAN as the dimension increases; see Table 1.
@@
Figure 4: Logistic dependence structure. Left: $d=10$; right: $d=50$. – First two margins of extremes in the test set and the different generative methods.
4.2.2 Hüsler–Reiss dependence structure
In this setup, the random vector $\bm{X}$ is generated using a $d$-dimensional multivariate Pareto distribution with Hüsler–Reiss dependence structure having a random parameter matrix and Pareto(2) margins. The parameter matrix generation and the random sampling are done using the R package graphicalExtremes [graphicalExtremes]. The family is parametrized by a variogram matrix $\Gamma=(\gamma_{ij})_{i,j=1}^{d}$ where $\gamma_{ij}=0$ means complete dependence and $\gamma_{ij}=+\infty$ means asymptotic independence.
We consider various scenarios in which we assess the performance of WA-GAN and compare it to GPGAN, whose performance in generating multivariate extremes, according to the previous section, is quite close to the results of our procedure. The models are trained on $n_{\text{train}}=10\,000$ observations and validated on $n_{\text{val}}=5\,000$ observations. The performance assessment is done on $n_{\text{test}}=20\,000$ points.
To illustrate the variety of tail strength dependence in the different scenarios, we plot the parameter matrix of each case as heatmaps in Appendix E in the supplement.
Results.
The setup and the meaning of the scores are exactly the same as explained above Table 1.
Dependence score Extremes score
Model $d=10$ $d=20$ $d=50$ $d=10$ $d=20$ $d=50$
WA-GAN $\bm{0.0145}$ $\bm{0.0176}$ $\bm{0.0123}$ $\bm{16.69}$ $\bm{9.512}$ $\bm{14.926}$
GPGAN 0.0629 0.1431 0.0989 20.979 11.406 21.067
Table 2: Hüsler–Reiss dependence structure. Dependence and extremes scores for different dimensions $d$. For each dimension and each score, the best score (i.e., the lowest) is indicated in bold.
WA-GAN remains highly competitive in this more complex scenario, both in terms of dependence and generated extremes. Even though the degrees of dependence vary considerably in this setting (Figure 5), WA-GAN is able to capture them quite accurately, even in high dimensions. For GPGAN, the situation is less favorable, as the deviation from the reference coefficients becomes quite pronounced when $d$ increases.
Figure 5: Hüsler–Reiss dependence structure. Left: $d=10$; right: $d=50$. – Extremal coefficients of order $k=2$ (blue) and $k=3$ (red).
We investigate WA-GAN’s sensitivity to the assumption that the angular measure is concentrated on ${\kern 0.0pt\Delta}^{\mathrm{o}}_{d-1}$. Indeed, by construction, the method does not allow the simulation of extremes associated with directions on the boundary of the simplex $\Delta_{d-1}$. Based on the parameter matrices shown in Figure 11 in Appendix E, we consider, for $d\in\{10,50\}$, two scenarios exhibiting high and low tail dependence between two components, and we compare the generated extremes to the test extremes in Figure 6. WA-GAN performs well when there is moderate to strong tail dependence but fails to generate extremes that are too close to the axes, corresponding to the low tail dependence scenario, especially if $d=50$. We discuss this issue in the next section under the asymptotic independence scenario and propose a possible improvement to the method.
Figure 6: Hüsler–Reiss dependence structure. Left: $d=10$; right: $d=50$. – Test extremes are in black while generated extremes by WA-GAN are in purple. For each dimension, the left plot corresponds to a high tail dependence scenario while the right plot corresponds to a low tail dependence scenario.
4.2.3 Gaussian dependence structure
In this setup, the random vector $\bm{X}$ is generated using a $d$-dimensional exchangeable Gaussian copula with correlation parameter $\rho\in[-1,1]$ and Pareto distributed marginals with parameter $\alpha>0$. It is easy to verify that this data-generating process does not satisfy our assumptions as Assumption 2 is not verified: all variables are asymptotically independent and the angular measure is concentrated on the vertices of the unit simplex.
We consider various scenarios in which we assess the performance of WA-GAN. We take $d\in\{10,20,50\}$ and a Kendall’s tau of $\tau\in\{1/4,1/2,3/4\}$, corresponding to the correlation parameters $\rho\in\{\sin(\pi\tau/2):\tau=0.25,0.5,0.75\}$. For the margins, we consider $\alpha=2$ in any scenario. The models are trained on $n_{\text{train}}=10\,000$ observations and validated on $n_{\text{val}}=5\,000$ observations. The performance assessment is done on $n_{\text{test}}=20\,000$ points. We do not report other methods’ performances, as our aim is mainly to see how we may adapt our methodology to be resilient against asymptotic independence in the data.
As our method relies on Aitchison coordinates for the open simplex, it is not able to produce new angles located exactly on the boundary of the simplex. However, one crucial observation is the following: lower tail dependence corresponds to heavier tails in the Aitchison coordinate space. This is illustrated in Figure 7 with $d=3$ and $n=1000$ data points sampled from a Dirichlet distribution with parameter vector $\alpha=(2,2,2)$ (top) and $\alpha=(0.1,0.1,0.1)$ (bottom). In each situation, we display the raw simplex data, the corresponding coordinates in the orthonormal basis introduced in Proposition 2, and the frequency histogram of the first coordinate values.
(a) Raw simplex data
(b) Aitchison coordinates
(c) Histogram of the first coordinate values
(d) Raw simplex data
(e) Aitchison coordinates
(f) Histogram of the first coordinate values
Figure 7: The Aitchison transform for Dirichlet distributed data in $d=3$ with parameter vector $\alpha=(2,2,2)$ (top row) $\alpha=(0.1,0.1,0.1)$ (bottom row).
Given this observation, we propose to allow for other latent distributions than the standard normal one as input to the generator of WA-GAN. Here, we consider the Student distribution with possible degrees of freedom equal to $1$, $2.5$, 55, or $\infty$, the latter value corresponding to the Gaussian distribution as before. The latent distribution choice is considered as a hyperparameter and is chosen randomly during the random search.
Results.
Table 3 reports the test scores, for the same two scores as used previously, of WA-GAN in each of the considered scenarios.
Dependence score Extremes score
$\tau$ $d=10$ $d=20$ $d=50$ $d=10$ $d=20$ $d=50$
$\tau=\frac{1}{4}$ 0.0185 0.0214 0.0259 36.939 063.489 45.041
$\tau=\frac{1}{2}$ 0.0335 0.0173 0.0166 53.859 117.450 71.547
$\tau=\frac{3}{4}$ 0.0135 0.0115 0.0108 39.144 099.651 74.948
Table 3: Gaussian dependence structure. Dependence and extremes scores for different dependence $\tau$ and dimensions $d$.
The dependence score values are quite similar to those obtained with the logistic model. However, regarding extreme values, the results are worse, as the method tends to produce overly dependent tail observations (Figure 8(a)). One possible explanation for the fact that the dependence score is not strongly affected is that, under asymptotic independence in the tail, the coefficients are very poorly estimated—even in the test set—so that the score values are not truly reliable in this case (Figure 8(b)).
(a) Gaussian dependence structure. First two margins of extremes in the test set and the generated ones using WA-GAN for $d=10$.
(b) Gaussian dependence structure. Extremal coefficients of order $k=2$ (blue) and $k=3$ (red). Solid black line is the diagonal. Dashed blue and red lines are the true values of the coefficients for $k=2$ and $k=3$ respectively.
Figure 8: Results for Gaussian dependence structure with $d=10$.
Despite the rather poor performance in generating asymptotically independent extremes, it is remarkable that, out of the nine possible scenarios, six of the validated architectures employed a Student rather than Gaussian latent distribution for the Aitchison coordinates. This suggests that allowing for heavier-tailed distributions may be beneficial when asymptotic independence is suspected. The selected degrees of freedom were four times $2.5$ and two times $5$, indicating that an excessively heavy-tailed distribution, with no finite mean, might be somewhat too extreme.
@@ 4.3 Financial data: daily returns of industry portfolios
We apply our methodology to analyze the tail behavior of the “value-averaged” daily returns of $d=30$ industry portfolios compiled and posted as part of the Kenneth French Data Library. The data in consideration span between 1950 and 2015 with n=16 694n=16\,694 observations. This data set is very widely used; for analyses in an extreme value context see e.g. [janssen2020] and [cooley2019]. Details about the portfolio constructions can be found online. [^1] https://mba.tuck.dartmouth.edu/pages/faculty/ken.french/Data_Library/det_30_ind_port.html Since we are interested in extreme losses we first multiply all returns by 1-1.
@@
To sample extremes from the joint distribution, we first compute the maximum likelihood estimators of the univariate GP parameters for each margin. Boxplots illustrating the values of those parameters are reported in Figure 12 in Appendix E. The estimated shape values clearly indicates the presence of heavy tails in the marginal distributions. The GP qq-plots of the marginal fits are to be found on Figures 13, 14 and 15 in Appendix E in the supplement. They indicate a relatively good fit, even though some extreme quantiles are larger than the ones associated with the GP model.
For comparison, we start by computing the test scores for WA-GAN, HTGAN and GPGAN. The training and validation procedures are exactly the same as described in Section 4.2. Here, we train on $n_{\text{train}}=7\,000$, we validate on $n_{\text{val}}=3\,000$ and compute the scores on $n_{\text{test}}=6\,694$ observations. We again consider $k=\sqrt{n_{\text{train}}}$ for WA-GAN. For HTGAN, we do not compute the scores on extremes because the marginal distributions are not known, see Appendix D. Table 4 shows that WA-GAN seems to be the best at capturing the dependence between extremes in the data while GPGAN has a better $2$-Wasserstein distance score.
@@
As already pointed out in the analysis of [janssen2020], the extreme losses of those portfolios can be clustered into different kind of industries. For example, the tobacco industry exhibits asymptotic independence to all other categories (e.g., to the textile industries) while the extreme losses of the business and IT-related industries are dependent. We illustrate the fact that WA-GAN also captures those phenomena in Figure 10(a) by displaying the two-dimensional projections of the estimated angular measure on those margins. We see that the generated “angles” associated with tobacco/textile portfolios are much more concentrated around the axes than for business/IT, showing that WA-GAN is able to capture quite reasonably this complex tail dependence in this setting.
As a further illustration we consider a portfolio obtained by combining the mines industry portfolio with the coal industry portfolio (with equal weight). It is intuitive and has been shown in [cooley2019, janssen2020] that the mines and coal industries exhibit asymptotic dependence. If one wants to sample the losses associated to the portfolio we propose, this tail dependence should be taken into account. Figure 10(b) represents the densities of simulated values of the extreme losses of the combined portfolio (i.e., both marginal losses exceeds a large threshold) based on WA-GAN (in black) and based on independent simulations from the marginal GP distributions (in blue). Not taking the tail dependence into account leads to underestimation of the risk associated with this portfolio.
(a) Two-dimensional projections of the generated “angles” from $\Phi$ by WA-GAN.
(b) Densities of extreme losses in the combined portfolio of mines and coals industries based on WA-GAN (black) and independently simulated margins (blue).
Figure 10: Simulation results for the financial dataset.
@@ 5 Conclusions
The method is tested on simulated extreme-value datasets of dimensions $d=10$, 2020, and $50$, with 10,00010{,}000 observations in the training set and 5,0005{,}000 in the validation set. Its performance is evaluated using two metrics: a dependence score based on extremal coefficients and a Wasserstein score illustrating the quality of the generated extremes. In both cases, WA-GAN performs reasonably well compared to other existing methods such as HTGAN and GPGAN. However, its performance may deteriorate under asymptotic independence, as it relies on the Aitchison space, where low tail dependence corresponds to heavier-tailed coordinates, while WA-GAN uses a Gaussian latent distribution. Possible improvements to address this limitation will be discussed in future work.
@@
An alternative approach would be to take a Bayesian perspective: choose a prior distribution for the marginal tail parameters and the angular measure (or any other means of describing the dependence structure) and then simulate from the predictive distribution given the data: first simulate random tail and parameters from the posterior distribution and then sample from the resulting multivariate generalized Pareto distribution. The advantage of such an approach is that the estimation uncertainty would be taken into account through the posterior distribution of the tail and dependence parameters. Drawbacks are that the construction of a suitable prior on the dependence structure may become quite delicate and the generation from the posterior distribution rather involved, especially in high dimensions.
Another important aspect, not addressed in this work but potentially relevant in applications, is the possible time-evolving nature of the extremal dependence structure of the data [drees2023time, deCarvalho2025techs], or its variation with respect to a latent variable. Covariates can be incorporated into (Wasserstein) GANs by using their conditional versions [mirza2014cGAN, zheng2020cWGANGP], which could facilitate the generation of new multivariate extremes for specific covariate values, like time, in a manner similar to what has been explored for image generation in [gauthiercGAN]. These directions are left for future work.
@@
The code is publicly available on the repository https://github.com/stephanelh98/extreme-WGAN. The data underlying the analysis in Section 4.3 is publicly available from the Kenneth R. French data library https://mba.tuck.dartmouth.edu/pages/faculty/ken.french/data_library.html.
@@ Appendix A Proof of Proposition 1
The exponent measure $\nu$ is connected to the angular measure $\Phi$ introduced in (2). More precisely [beirlant2004, Chapter 8], for any bounded, measurable function $f:\mathbb{E}\rightarrow\mathbb{R}$ that is zero in a neighborhood of the origin, we have
@@
where $Y$ is unit-Pareto distributed and independent of $\bm{\Theta}\sim\Phi$. Hence, by (26),
@@
From the local uniformity of the convergence in point (ii) of Remark 1, guaranteed by Assumption 3, and the fact that $\mathsf{P}(Y_{j}>0)=1$ for any $j\in\{1,\ldots,d\}$ by the previous computations and Assumption 2, the extended continuous mapping theorem [van1996weak, Theorem 1.11.1] applies, yielding
@@
This concludes the proof of (7) and (8). Eq. (9) is a direct consequence of (7) as
@@ Appendix B The Aitchison simplex
The open simplex ${\kern 0.0pt\Delta}^{\mathrm{o}}_{d-1}$ is equipped with an inner product space structure with the following operations [aitchison1982]: for $\bm{v}=(v_{1},\ldots,v_{d}),\bm{w}=(w_{1},\ldots,w_{d})\in{\kern 0.0pt\Delta}^{\mathrm{o}}_{d-1}$ and $\alpha\in\mathbb{R}$, define the following operations:
$\bm{v}\oplus\bm{w}$ $\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\left(\frac{v_{j}w_{j}}{\sum_{i=1}^{d}v_{i}w_{i}}\right)_{j=1}^{d}$ (addition),
$\alpha\odot\bm{v}$ $\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\left(\frac{v_{j}^{\alpha}}{\sum_{i=1}^{d}v_{i}^{\alpha}}\right)_{j=1}^{d}$ (scalar multiplication),
$\left\langle\bm{v},\bm{w}\right\rangle_{A}$ $\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\sum_{i=1}^{d}\log\left(\frac{v_{i}}{g(\bm{v})}\right)\log\left(\frac{w_{i}}{g(\bm{w})}\right)$ (Aitchison inner product),
where $g:\bm{u}\in{\kern 0.0pt\Delta}^{\mathrm{o}}_{d-1}\mapsto g(\bm{u})\stackrel{{\scriptstyle\mathrm{def}}}{{=}}(\prod_{i=1}^{d}u_{i})^{1/d}$ is the geometric mean. It is an easy exercise to verify that $({\kern 0.0pt\Delta}^{\mathrm{o}}_{d-1},\oplus,\odot)$ forms a real vector space with zero element $\bm{0}_{A}=\bm{1}_{d}/d\in{\kern 0.0pt\Delta}^{\mathrm{o}}_{d-1}$, where $\bm{1}_{d}=(1,\ldots,1)\in\mathbb{R}^{d}$, and that $\left\langle\cdot,\cdot\right\rangle_{A}$ defines an inner product on it. The open simplex equipped with this structure is often called the Aitchison simplex.
@@
where $\mathbb{H}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\{\bm{x}\in\mathbb{R}^{d}:\langle\bm{x},\bm{1}_{d}\rangle=0\}$ with $\langle\cdot,\cdot\rangle$ the standard inner product in $\mathbb{R}^{d}$, then
@@
$\bigl\{\bm{x}_{i}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}(0,\ldots,0,1,-1,0,\ldots,0)\in\mathbb{H}:i=1,\ldots,d-1\bigr\},$